# Sign-In to the Lottery: Reparameterizing Sparse Training From Scratch

### v1

@@ 4 Sign-In: Learning Sign Flips
This holds for general continuously differentiable objective function $L:\mathbb{R}^{m}\rightarrow\mathbb{R}$. To induce helpful sign flips, we have to deal with two factors: stochastic noise (Pesme et al., 2021) and weight decay (Jacobs & Burkholz, 2025). Both factors have the same effect on the training dynamics: They induce sparsity by shrinking $\beta$. Yet, if $\beta=0$, we cannot sign flip due to the $\sqrt{\theta_{t}^{2}}$ in the Riemannian gradient flow Eq. (1). We address this issue with Algorithm 1 by resetting the inner scaling $\beta$ without changing the weights of the network every $p$ epochs up to epoch $T_{2}$. In all our experiments, we choose the same scaling of $\beta=1$ for all parameters, as this leads to stable results. We employ an analytic solution to apply the rescaling on both $m$ and $w$, which is given in Appendix B. This gives the weights more opportunities to align their signs correctly, as we prove for a simple example in Section 5.
Computational and memory cost
Sign-In doubles the amount of parameters relative to sparse training. This is also the case for continuous sparsification methods that utilize the implicit bias of $m\odot w$ towards sparsity, like spred (Ziyin & Wang, 2023) and PILoT (Jacobs & Burkholz, 2025). Nevertheless, according to (Ziyin & Wang, 2023), the training time of a ResNet50 with $m\odot w$ parameterization on ImageNet increases roughly by 5% only and the memory cost is negligible if the batch size is larger than $50$. Furthermore, at inference, we would return to the representation $\theta$ by merging $\theta=m\odot w$. Therefore, Sign-In is able to improve sparse training with minimal overhead.
Algorithm 1 Sign-In
  for $i=1$ to $T$ do
     if $i$ mod $p==0$ and $i<T_{2}$ then
     end if
  end for
  Return Model parameters $\theta_{T}=S\odot m_{T}\odot w_{T}$
5 Sign Learning in Theory
Gadhikar & Burkholz (2024a) proposed a simple testbed to compare two iterative pruning algorithms, IMP and LRR. LRR’s superior performance is explained by its ability to flip the sign of a critical parameter during a dense training phase. As we show, the same mechanism distinguishes sparse training from scratch with random initialization and dense training. The next section is dedicated to the question how Sign-In changes this picture. Remarkably, it provably resolves another case than dense (overparameterized) training. In combination with dense training, it can flip even all relevant signs and solve the learning problem (see Figure 2).
Single neuron setting
Let $\{\mathbf{z}_{i}\}_{i=1}^{n}$ be a data set of i.i.d $\mathcal{N}(\mathbf{0},\mathbf{I})$ random variables. Consider a two-layer neural network consisting of a single hidden neuron $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that $f(\mathbf{z}\mid a,\mathbf{w}):=a\sigma\left(\mathbf{w}^{T}\mathbf{z}\right)$, where $\sigma(\cdot):=\max\{\cdot,0\}$ denotes the Rectified Linear Unit (ReLU). The labels $\{y_{i}\}_{i=1}^{n}$ are generated from a sparse teacher $y_{i}=\tilde{f}\left(\mathbf{z}_{i}\mid\tilde{a},\tilde{\mathbf{w}}\right)$ with $\tilde{a}>0$ and $\tilde{\mathbf{w}}=(1/\tilde{a},0,0,..,0)$. On this data, we train a student two-layer neural network with gradient flow and randomly balanced parameter initialization, i.e., $a^{2}_{\text{in}}=\left\lVert\mathbf{w_{\text{in}}}\right\rVert^{2}_{2}$. The objective is to minimize the mean squared error (MSE), which is given by $L(a,\mathbf{w}):=\frac{1}{2n}\sum_{i=1}^{n}\left(f\left(\mathbf{z}_{i}\mid a,\mathbf{w}\right)-y_{i}\right)^{2}$. The training dynamics with loss function $L$ is described by the gradient flow:
@@ 5 Sign Learning in Theory
(Theorem 2.1 (Gadhikar & Burkholz, 2024a)) In the one-dimensional single neuron setting, Eq. (2) with $d=1$, the student can recover the ground truth given sufficiently many samples, if $a_{in}>0$ and $w_{1,in}>0$. In all other cases ($a_{in}>0$, $w_{1,in}<0$), ($a_{in}<0$, $w_{in}>0$) and ($a_{in}<0$, $w_{1,in}<0$), it will not attain the correct solution.
[…]
Sign-In reparameterization
The Sign-In reparameterization, which we have introduced in Section 4, can resolve the additional case $a_{in}<0,w_{1,in}>0$ when $d=1$. Note that both $a$ and $w$ are reparameterized with different inner scalings denoted by $\beta_{1}$ and $\beta_{2}$. The corresponding Riemannian gradient flow in the original parameters $a$ and $w_{1}$ is:
@@
To illustrate Theorem 5.2, we simulate the toy setup for both the standard gradient flow and the reparameterized gradient flow. Figure 3 shows that the standard gradient flow converges only when the signs are correctly initialized, whereas the reparameterized model also succeeds if $a_{in}<0,w_{1,in}>0$. Other initializations lead the student to converge to the origin and thus a function that is constant at 00, i.e. training fails.
Figure 3: Student-teacher setup training one neuron For a single input $d=1$, when $w_{1}\leq 0$ both methods fail to reach the ground truth. If $w_{1}>0$, the Sign-In gradient flow succeeds, whereas gradient flow fails in this case when additionally $a<0$.
@@
Multi-dimensional input
For multi-dimensional input ($d>1$), the outer layer parameter $a$ cannot sign flip like in case of $d=1$ (Lemma 2.2 (Gadhikar & Burkholz, 2024a)). However, we have shown that when the rescaling of $a$ is larger then that of $w_{1}$, i.e., $\beta_{1}>\beta_{2}>0$, the sign of $a$ is recovered, if $w_{1}$ initialized with the correct sign. This is impractical for larger neural networks with more layers, since we have to choose a different scaling $\beta$ for all layers. Nonetheless, for a multi-dimensional input layer, different rescaling is not necessary, as we verify empirically next.
@@
A potential explanation is that a finite learning rate $\eta>0$ induces overshooting, leading to sign flips. However, gradient flow with the standard parameterization does not recover the correct sign of $a$ due to this overshooting. Accordingly, the success of Sign-In cannot be attributed entirely to this factor. The key factor is that the balance between $a$ and $||\mathbf{w}||_{L_{2}}$ is broken. Intuitively, for $a_{t}$ to sign flip, it needs to move faster than $||\mathbf{w}_{t}||_{L_{2}}$ as in the one-dimensional case (see Figure 3). We observe that $|a_{in}|^{2}=||\mathbf{w}_{in}||^{2}_{L_{2}}\geq w_{k,in}^{2}$ for all $k\in[d]$. Therefore, the reparameterization induces a relative speed up for $a$ compared to $||\mathbf{w}||_{L_{2}}$, i.e.,
$\sqrt{a^{2}_{in}+\beta}=\sqrt{||\mathbf{w}_{in}||^{2}_{L_{2}}+\beta}\geq\frac{1}{n}\sum_{k=1}^{n}\sqrt{w_{k,in}^{2}+\beta}$ (4)
@@
No replacement for overparameterization
While it is remarkable that the combination of Sign-In and overparameterized training can solve our theoretical toy problem, ideally, we would like to facilitate training from scratch ($d=1$) without utilizing any dense training. Sign-In is not able to achieve this without dense training but does there exist another reparameterization that could? Unfortunately, the answer is negative, as we establish next.
@@
Multiple neurons
Do our insights still hold if we increase the number of neurons? To verify the broader validity of our claims, we consider a neural network $f_{2}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ given by $f_{2}(\mathbf{z}\mid\mathbf{a},\mathbf{W}):=\sum_{i=1}^{k}a_{i}\sigma\left(\mathbf{w}_{i}^{T}z\right)$. The data is generated from a similarly structured teacher network $\tilde{f}_{2}$ with $k=3$ neurons and input dimension $d=2$. Furthermore, $\tilde{a}_{i}$ are i.i.d. Rademacher random variables, i.e. $\mathbf{\tilde{a}}\sim Uni(\{-1,1\})$, and $\tilde{w}_{i,j}\sim N(0,1)$ i.i.d. for all $i,j\in[k,d]$ and normalized over the input dimension. We initialize the dense network with $k=20$ and the so-called COB initialization based on the rich regime in (Chizat et al., 2020). All weights are initialized with $N(0,1/d)$ and we ensure that $a_{i}=-a_{i+10}$ for $i\in[10]$. In the sparse case, we learn $k=3$ neurons. A difficult sign initialization is chosen as $a_{i}>0$ for all $i\in[3]$. The learning rate used is $\eta=1$ for the sparse setup and $\eta=2.15$ for the dense setup. The scaling for Sign-In is $\beta=2$. Figure 4 illustrates each neuron by a vector $|a_{i}|\mathbf{w}_{i}$. The teacher neurons are vectors touching the unit circle and student neurons are denoted with markers. Sparse initialization with good alignment and the dense case are in Appendix C. Also in these cases, Sign-In usually resolves the failure cases of gradient flow.
(a) Gradient flow.
(b) Sign-In gradient flow.
Figure 4: Representation of a two layer student-teacher neural network with multiple neurons inspired by Chizat et al. (2020). (a) Sparse training from scratch fails with bad signs. (b) Sign-In enables learning the representation with bad signs in Fig. 4(b).
@@ 6 Experiments
We show that Sign-In provably improves PaI in a small scale example (Theorem 5.2), now we have to verify that its merits also translate to larger scale settings. While our main focus is on PaI and thus training sparse networks from scratch, our theory implies that Sign-In could also boost dense-to-sparse training performance, as their combination can resolve cases where both alone would fail. However, we do not expect training from scratch with Sign-In to outperform training starting from well aligned signs, as Sign-In acts orthogonally to dense-to-sparse training in our theoretical setting. Sign-In is expected to facilitate different sign flips.
Experimental Setup
We perform experiments on standard vision benchmarks including CIFAR10, CIFAR100 (Krizhevsky et al., 2009), and ImageNet (Deng et al., 2009) training a ResNet20, ResNet18, and ResNet50 model, respectively. For architectural variation, we also train a vision transformer DeiT Small (Touvron et al., 2021) on ImageNet. All training details are provided in Table 6 in Appendix D.
@@
Scaling reset
An integral part of Sign-In is that we dynamically reset the internal scaling to combat the implicit bias towards sparsity of the parameterization. To ablate its contribution, we report results for Sign-In with and without resetting the scaling for a random mask trained from scratch in Table 11 in the appendix. Indeed, without the scaling reset, the performance gain is less significant. This confirms that resetting can mitigate the effects of stochastic noise and weight decay and enable better sign flips.
Sign-In improves sparse training
Table 3 reports results for sparse training with a random mask. The layerwise sparsity ratios are chosen as balanced such that each layer has the same number of nonzero parameters (Gadhikar et al., 2023). Sparse training with Sign-In as per Algorithm 1 consistently outperforms standard training. A similar improvement is observed for other PaI methods in Table 9 in Appendix G. We apply Sign-In also to different sparse masks identified with PaI or dense-to-sparse methods as reported in Table 10 in Appendix G. While layerwise sparsity ratios determine the overall trainability of a sparse network, Sign-In improves performance in better trainable settings.
Vision transformer with Sign-In
Sign-In is also able to improve sparse training of a randomly masked vision transformer. A DeiT Small architecture (trained for $300$ epochs from scratch) gains more than $1\%$ in performance with Sign-In, as reported in Table 4.
Table 4: DeiT Small on ImageNet, trained from scratch.
Method $s=0.5$
Random $69.31(\pm 0.7)$
Random + Sign-In (ours) $\mathbf{70.62(\pm 0.4)}$
Improved sign alignment
Figure 5(a) validates the Sign-In mechanism by tracking the percentage of sign flips during training for a $90\%$ sparse ResNet50 on ImageNet. Across the full duration of training, Sign-In enables strictly more sign flips than sparse training. This improves sign alignment, as reflected in the improved performance and the flatter minima, which are obtained by Sign-In (see Figure 5(b)). Since Sign-In employs an alternate mechanism for sign flipping, it cannot learn stable signs early like dense-to-sparse methods. Yet, it enables more sign flips throughout training as compared to standard sparse training. This aids alignment by better exploration of the parameter space.
@@
Sign-In can also improve dense-to-sparse training
Only a combination of Sign-In and overparameterization is able to solve all sign flip challenges in our theoretical setting, as also confirmed by numerical experiments, which are provided in Table 8 in Appendix E. Therefore, we expect the combination of dense-to-sparse methods with Sign-In to achieve further gains in performance. Table 5 confirms that AC/DC improves with Sign-In. However, the improvement is less significant compared to PaI.
Implications of theory
The surprising success of Sign-In combined with overparameterization in our theoretical setting relies on the last layer $a$ learning at a higher speed than the first layer $w_{1}$. This enables $a$ to flip its sign if necessary, which is impossible in the same setting without Sign-In. $w_{1}$ can follow once the sign of $a$ is correctly learned. We conjecture that weights in layers close to the input would have a harder time flipping in deep neural networks consisting of more layers and neurons, as other layers on top slow its learning down. In line with this narrative, Sign-In promotes more sign flips in layers closer to the output, as shown in Figure 8 in the appendix. To understand how these sign flips relate to generalization, we must analyze more complex scenarios in future and see whether more effective sign learning could further boost dense-to-sparse training. The primary focus of this work has been to improve sparse training from scratch and Sign-In overcomes one of its major limitations regarding learning parameter signs.
Table 5: Sign-In improves AC/DC for a ResNet50 trained on ImageNet.
Method $s=0.8$ $s=0.9$ $s=0.95$
AC/DC $75.83(\pm 0.02)$ $74.75(\pm 0.02)$ $72.59(\pm 0.11)$
+ Sign-In
@@ 7 Conclusion
Parameter signs that are aligned with a given mask define an initialization that is sufficient to train a sparse network from scratch. We have empirically verified this conjecture for masks that result from state-of-the-art dense-to-sparse training algorithms. This is enabled by access to well aligned signs, which are learned early during dense-to-sparse training. Regardless of the sparse mask being task specific or not, learning highly performant parameter signs is a great challenge for PaI. To aid parameter sign flipping, we have proposed to reparameterize a neural network by Sign-In and have demonstrated its effectiveness in a diverse range of experiments as well as a theoretically tractable setting. For a single neuron network, we have proven that Sign-In resolves a case that is complementary to overparameterized training. Their combination, overparameterization with Sign-In, even recovers the ground truth in all possible scenarios. Ideally, we would like to make dense (overparameterized) training phases obsolete in pruning methods to save computational resources. Although, as we have shown, a reparameterization alone will not be able to achieve this. Closing the gap between training sparse, randomly initialized networks from scratch and dense-to-sparse training remains a hard open problem. Our work brings us closer to this goal with two key insights: a) sign alignment is enough for training a sparse initialized network and b) reparameterization can improve sign alignment.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v2

@@ 4 Sign-In: Learning Sign Flips
To improve the learning of task relevant sign flips, we introduce a reparameterization. Every weight parameter $\theta$ is replaced by the product of two parameters $m\odot w$, where $\odot$ denotes a pointwise multiplication (Hadamard product). This introduces an additional degree of freedom $\beta$, which is determined by the initialization and captured by $m^{2}-w^{2}=\beta\mathbf{1}$, where $\beta>0$ and $\mathbf{1}=(1,...,1)$. According to Lemma 4.8 (Li et al., 2022), the reparameterization induces a Riemannian gradient flow in the original parameter $\theta$:
This holds for general continuously differentiable objective function $L:\mathbb{R}^{m}\rightarrow\mathbb{R}$. To induce helpful sign flips, we have to deal with two factors: stochastic noise (Pesme et al., 2021) and weight decay (Jacobs & Burkholz, 2025). Both factors have the same effect on the training dynamics: They induce sparsity by shrinking $\beta$. Yet, if $\beta=0$, we cannot sign flip due to the $\sqrt{\theta_{t}^{2}}$ in the Riemannian gradient flow Eq. (1). We address this issue with Algorithm 1 in Appendix˜B by resetting the inner scaling $\beta$ without changing the weights of the network every $p$ epochs up to epoch $T_{2}$. In all our experiments, we choose the same scaling of $\beta=1$ for all parameters, as this leads to stable results. We employ an analytic solution to apply the rescaling on both $m$ and $w$, as given in Appendix C. This gives the weights more opportunities to align their signs correctly, as proven for a simple example in Section 5.
Computational and memory cost Sign-In doubles the amount of parameters relative to sparse training. This is also the case for continuous sparsification methods that utilize the implicit bias of $m\odot w$ towards sparsity, like spred (Ziyin & Wang, 2023) and PILoT (Jacobs & Burkholz, 2025). Nevertheless, according to Ziyin & Wang (2023), the training time of a ResNet50 with $m\odot w$ parameterization on ImageNet increases roughly by 5% only and the memory cost is negligible if the batch size is larger than $50$. Furthermore, at inference, we would return to the representation $\theta$ by merging $\theta=m\odot w$. Therefore, Sign-In is able to improve sparse training with minimal overhead as also highlighted by a FLOP comparison in Table˜15.
5 Sign Learning in Theory
Gadhikar & Burkholz (2024) proposed a simple testbed to compare two iterative pruning algorithms, IMP and LRR. LRR’s superior performance is explained by its ability to flip the sign of a critical parameter during a dense training phase. As we show, the same mechanism distinguishes sparse training from scratch with random initialization and dense training. The next section is dedicated to the question how Sign-In changes this picture. Remarkably, it provably resolves another case than dense (overparameterized) training. In combination with dense training, it can flip even all relevant signs and solve the learning problem (see Figure 2).
SparsewwaaOverparamwwaaSparse+ Sign-InwwaaOverparam+ Sign-Inwwaa
@@ 5 Sign Learning in Theory
(Theorem 2.1 (Gadhikar & Burkholz, 2024)) In the one-dimensional single neuron setting, Eq. (2) with $d=1$, the student can recover the ground truth given sufficiently many samples, if $a_{in}>0$ and $w_{1,in}>0$. In all other cases ($a_{in}>0$, $w_{1,in}<0$), ($a_{in}<0$, $w_{in}>0$) and ($a_{in}<0$, $w_{1,in}<0$), it will not attain the correct solution.
Sign-In reparameterization The Sign-In reparameterization, which we have introduced in Section 4, can resolve the additional case $a_{in}<0,w_{1,in}>0$ when $d=1$. Note that both $a$ and $w$ are reparameterized with different inner scalings denoted by $\beta_{1}$ and $\beta_{2}$. The corresponding Riemannian gradient flow in the original parameters $a$ and $w_{1}$ is:
@@
To illustrate Theorem 5.2, we simulate the toy setup for both the standard gradient flow and the reparameterized gradient flow. Figure 3(a) shows that the standard gradient flow converges only when the signs are correctly initialized, whereas the reparameterized model also succeeds if $a_{in}<0,w_{1,in}>0$. Other initializations lead the student to converge to the origin and thus a function that is constant at 0, i.e. training fails.
@@
Multi-dimensional input For multi-dimensional input ($d>1$), the outer layer parameter $a$ cannot sign flip like in case of $d=1$ (Lemma 2.2 (Gadhikar & Burkholz, 2024)). However, we have shown that when the rescaling of $a$ is larger then that of $w_{1}$, i.e., $\beta_{1}>\beta_{2}>0$, the sign of $a$ is recovered, if $w_{1}$ initialized with the correct sign. This is impractical for larger neural networks with more layers, since we have to choose a different scaling $\beta$ for all layers. Nonetheless, for a multi-dimensional input layer, different rescaling is not necessary, as we verify empirically next.
@@
[…]
No replacement for overparameterization While it is remarkable that the combination of Sign-In and overparameterized training can solve our theoretical toy problem, ideally, we would like to facilitate training from scratch ($d=1$) without utilizing any dense training. Sign-In is not able to achieve this without dense training but does there exist another reparameterization that could? Unfortunately, the answer is negative, as we establish next.
@@
Multiple neurons Do our insights still hold if we increase the number of neurons? To verify the broader validity of our claims, we consider a neural network $f_{2}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ given by $f_{2}(\mathbf{z}\mid\mathbf{a},\mathbf{W}):=\sum_{i=1}^{k}a_{i}\sigma\left(\mathbf{w}_{i}^{T}z\right)$. The data is generated from a similarly structured teacher network $\tilde{f}_{2}$ with $k=3$ neurons and input dimension $d=2$. The weights are initialized and trained as outlined in Appendix˜D. A difficult sign initialization is chosen as $a_{i}>0$ for all $i\in[3]$. Figure 3 illustrates each neuron by a vector $|a_{i}|\mathbf{w}_{i}$. The teacher neurons are vectors touching the unit circle and student neurons are denoted with markers.
(a) Single neuron.
(b) Gradient flow.
(c) Sign-In.
Figure 3: Student-teacher setup (a) Training one neuron for a single input $d=1$, when $w_{1}\leq 0$ both methods fail to reach the ground truth. If $w_{1}>0$, the Sign-In gradient flow succeeds, whereas gradient flow fails in this case when additionally $a<0$. (b) Representation of a two layer student-teacher neural network with multiple neurons inspired by Chizat et al. (2020) for sparse training from scratch fails with bad signs. (c) Sign-In enables learning the representation with bad signs.
@@ 6 Experiments
We show that Sign-In provably improves PaI in a small scale example (Theorem˜5.2), now we have to verify that its merits also translate to larger scale settings. While our main focus is on PaI and thus training sparse networks from scratch, our theory implies that Sign-In could also boost dense-to-sparse training performance, as their combination can resolve cases where either would fail. However, we do not expect training from scratch with Sign-In to outperform training starting from well aligned signs, as Sign-In acts orthogonally to dense-to-sparse training in our theoretical setting. Sign-In is expected to facilitate different sign flips.
Experimental Setup We perform experiments on standard vision benchmarks including CIFAR10, CIFAR100 (Krizhevsky et al., 2009), and ImageNet (Deng et al., 2009) training a ResNet20, ResNet18, and ResNet50 model, respectively. For architectural variation, we also train a vision transformer DeiT Small (Touvron et al., 2021) on ImageNet. All training details are provided in Table˜4 in Appendix E.
@@
Scaling reset An integral part of Sign-In is that we dynamically reset the internal scaling to combat the implicit bias towards sparsity of the parameterization. To ablate its contribution, we report results for Sign-In with and without resetting the scaling for a random mask in Table˜14 in the appendix. Indeed, without the scaling reset, the performance gain is less significant. This confirms that resetting mitigates the effect of stochastic noise and weight decay, enabling better sign flips.
Sign-In improves sparse training Table 3 reports results for sparse training with a random mask. The layerwise sparsity ratios are chosen as balanced such that each layer has the same number of nonzero parameters (Gadhikar et al., 2023). Sparse training with Sign-In as per Algorithm˜1 consistently outperforms standard training, substantiating our theoretical results. A similar improvement is observed for other PaI methods in Table˜7 in Appendix H. We apply Sign-In also to different sparse masks identified with PaI or dense-to-sparse methods as reported in Table˜8 in Appendix˜H. While layerwise sparsity ratios determine the overall trainability of a sparse network, Sign-In improves performance in better trainable settings.
Vision transformer with Sign-In Sign-In is also able to improve sparse training of a randomly masked vision transformer. A DeiT Small architecture (trained for $300$ epochs from scratch) gains more than $1\%$ in performance with Sign-In, as reported in Table˜10, Appendix˜J.
Improved sign alignment Figure 4(a) validates the Sign-In mechanism by tracking the percentage of sign flips during training for a $90\%$ sparse ResNet50 on ImageNet. Across the full duration of training, Sign-In enables strictly more sign flips than sparse training. This improves sign alignment, as reflected in the improved performance and the flatter minima (see Figure˜4(b)). Since Sign-In employs an alternate mechanism for sign flipping, it cannot learn stable signs early like dense-to-sparse methods. Instead, it enables more sign flips throughout training as compared to standard sparse training, to aid alignment by better exploration of the parameter space.
Sign-In improves other sparse training methods Only a combination of Sign-In and overparameterization is able to solve all sign flip cases in our theoretical setting, as also confirmed by numerical experiments in Table˜6, Appendix˜F. Therefore, we expect the combination of dense-to-sparse or dynamic sparse training methods with Sign-In to achieve further gains in performance. Appendix˜K shows that both AC/DC and RiGL are improved with Sign-In.
@@
Implications of theory The surprising success of Sign-In combined with overparameterization in our theoretical setting relies on the last layer $a$ learning at a higher speed than the first layer $w_{1}$. This enables $a$ to flip its sign if necessary, which is impossible in the same setting without Sign-In. $w_{1}$ can follow once the sign of $a$ is correctly learned. We conjecture that weights in layers close to the input would have a harder time flipping in deep neural networks consisting of more layers and neurons, as other layers on top slow its learning down. In line with this narrative, Sign-In promotes more sign flips in layers closer to the output, as shown in Figure˜7 in the appendix. To understand how these sign flips relate to generalization, we must analyze more complex scenarios in future and see whether more effective sign learning could further boost dense-to-sparse training. The primary focus of this work has been to improve sparse training from scratch and Sign-In overcomes one of its major limitations regarding learning parameter signs.
@@ 7 Conclusion
Parameter signs that are aligned with a given mask are sufficient to train a sparse network from scratch. We have empirically verified this conjecture for masks that result from dense-to-sparse training algorithms. Regardless of the sparse mask, learning highly performant parameter signs is a great challenge for PaI. To aid parameter sign flipping, we have proposed to reparameterize a neural network by Sign-In. For a single neuron network, we have proven with theory from Riemannian gradient flows and dynamical systems that Sign-In resolves a case that is complementary to overparameterized training which translates to largescale performance. Ideally, we would like to make dense training phases obsolete to save computational resources. Although, as we have shown, a reparameterization alone will not be able to achieve this. Closing the gap between training sparse, randomly initialized networks from scratch and dense-to-sparse training remains a hard open problem. Our work brings us closer to this goal with two key insights: a) sign alignment is enough for training a sparse initialized network and b) reparameterization can improve sign alignment.
8 Acknowledgements
We gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC) Jülich Supercomputing Centre (2021). We are also grateful for funding from the European Research Council (ERC) under the Horizon Europe Framework Programme (HORIZON) for proposal number 101116395 SPARSE-ML.